### v3

# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

Abstract
We derive formulae that resum, at a given order in the limit of small frequency, the infinite series of Post-Minkowkian (small gravitational coupling) or Post-Newtonian (small velocities) corrections to the gravitational waveform produced by particles moving along a general (open or closed) trajectory in Schwarzschild geometry. Specifying to the case of circular orbits, we compute the wave form and the energy flux to order 30PN, and compare it against the available results in the literature. The results are based on a novel hypergeometric representation of the solutions of the Heun equation (and its confluence), that leads to a mathematical proof of the Heun connection formula.
1 Introduction
Recent advances in the sensitivity of experimental apparatus for detecting gravitational waves (GWs) necessitate increasingly precise theoretical computations. The next generation of space-based experiments, such as the Laser Interferometer Space Antenna (LISA), promises to revolutionize the GW detection by significantly enhancing sensitivity and exploring entirely new wavelength ranges compared to ground-based interferometers LISA:2022kgy . LISA will detect long-wavelength GWs originating from sources like supermassive black hole mergers, extreme mass-ratio inspirals (EMRIs), galactic binaries, and the Stochastic Gravitational Wave Background—sources largely inaccessible to ground-based observatories.
The gravitational signals from these sources are diverse. For instance, GWs from supermassive black hole mergers will produce strong signals with large signal-to-noise ratios, while those from EMRIs, though weaker, will carry exceptionally detailed information about the gravitational sources. Indeed, EMRIs waveforms, spanning thousands of cycles during their inspiral phase, provide a unique opportunity to probe fundamental questions about the nature of gravitational sources, including their compactness, the presence or absence of horizons, light rings, echoes Cardoso:2019rvt , and the validity of Einstein’s theory of gravity and its possible extensions Gair:2012nm ; Yunes:2013dva .
Two main analytical tools have been developed for studying the problem we just described: the multipole expansion Thorne:1980ru ; Damour:1982wm ; Blanchet:2013haa and perturbation theory Teukolsky:1973ha ; Poisson:1993vp ; Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk . The latter, which applies to EMRI systems, is the focus of this paper. In this framework, GWs are computed by solving the Einstein equations to linear order in the perturbations induced by the motion of a lighter body orbiting around a very massive black hole (BH). The radial and angular propagations of the waves are governed by ordinary differential equations of the confluent Heun type. Analytic solutions of these equations are not known, but it is possible to write them as an expansion in the Newton constant (Post-Minkowskian, or PM, approximation) or in the orbital velocity (Post-Newtonian, or PN, approximation).
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in Fujita:2012cm . The expansion is known to be converging slowly Cutler:1993vq ; Poisson:1993vp due to the presence of logarithmic terms, leading to an infinite tower of tails and tails of tails non-local interactions Damour:2008gu ; Damour:2009kr ; Buonanno:2009zt ; Damour:2012ky . In Aminov:2020yma ; Bianchi:2021xpr ; Bianchi:2021mft ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms were proposed in terms of the instanton partition function of a SU(2)2 quiver gauge theory. Following this path the infinite tower of logarithmic terms have been resummed into exponentials.
More recently, multipole Bini:2023fiz ; Brandhuber:2023hhy ; DeAngelis:2023lvf ; Brandhuber:2023hhl ; Herderschee:2023fxh ; Georgoudis:2023eke ; Georgoudis:2023lgf ; Georgoudis:2023ozp ; Elkhidir:2023dco ; Bini:2024ijq ; Bini:2024rsy ; Georgoudis:2024pdz ; Alessio:2024onn ; Brunello:2024ibk and perturbation theory Fucito:2024wlg methods have been applied to the study of scattering processes (open trajectories) in gravity Kovacs:1977uw ; Kovacs:1978eu , a field where we can largely benefit from the impressive developments and refinements of scattering amplitude techniques in the last few years (see DiVecchia:2023frv for a review). These systems provide a richer set up, where the couplings of the PM and PN expansions are no longer linked to each other. For example, in the scattering amplitude language, the PM expansion corresponds to the expansion in loops. At a given order in the coupling, the amplitude is exact in the velocity, i.e. it resums the infinite PN series. To reproduce this result in perturbation theory, one has to resum the double expansion in one of the two variables.
For simplicity we focus on the geodesic motion in a Schwarzschild geometry and parametrize the mass, wavefrequency and radial distance by the two dimensionless combinations $z=2M/r$ and $y=2i\omega r$, assuming $x=4{\rm i}\omega M$ is small. We will show that, at order $x^{k}$, the dependence of the waveform on the variable $y$ or $z$ can be resummed and written in terms of a single hypergeometric function and its derivative, with coefficients given by two polynomials of degree $k$, that can be determined recursively. The two limits where $x$ is taking to be small keeping finite $y$ or $z$ lead to two alternative hypergeometric representations of the confluent Heun function, describing the PM and PN expansions respectively.
We observe that, the hypergeometric representation of the confluent Heun function we derived here is similar in spirit to the well studied MST solution Mano:1996vt . The crucial difference between the two, is that while the MST ansatz involves an infinite sum of hypergeometric functions, our solution is written in terms of a single hypergeometric function and its first derivative.
This leads to a mathematical proof of the Heun connection formulae originally found in Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten . The whole non-triviality of the Heun equation (and its confluences) is codified in a single function, $a(\ell)$, the Seiberg-Witten (SW) period of the underlying ${\cal N}=2$ supersymmetric gauge theory (see also Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe ). This function is related to the so called "renormalized angular momentum" in the MST context Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk or "anomalous dimension" in the effective field theory context Goldberger:2009qd ; Trestini:2023wwg ; Almeida:2021jyt ; Edison:2023qvg . The Heun connection formula has been recently applied in a variety of physical contexts, including the study of QNMs for gravity backgrounds (BHs, fuzzballs, D-branes), correlators in thermal field theories, holography, cosmological perturbations and scattering amplitudes in gravity Bianchi:2021xpr ; Bianchi:2021mft ; Fioravanti:2021dce ; Bianchi:2022wku ; Dodelson:2022eiz ; Dodelson:2022yvn ; Imaizumi:2022qbi ; Fioravanti:2022bqf ; Imaizumi:2022dgj ; Bianchi:2022qph ; Dodelson:2023vrw ; Bianchi:2023rlt ; Bianchi:2023sfs ; Giusto:2023awo ; Aminov:2023jve ; BarraganAmado:2023apy ; Lei:2023mqx ; BarraganAmado:2023wxt ; Fucito:2023afe ; Bautista:2023sdf ; Ge:2024jdx ; Cipriani:2024ygw ; Arnaudo:2024rhv ; Bianchi:2024vmi ; Bianchi:2024mlq ; DiRusso:2024hmd ; Bianchi:2024rod ; Aminov:2024aan ; BarraganAmado:2024tfu ; Bautista:2024emt ; Liu:2024eut ; Silva:2025khf ; Aprile:2025hlt ; Bhatta:2025kil .
To keep the material self-contained, in this paper we present a derivation of the Heun solutions from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. Finally, we remark, that even if the main motivation for resuming PM and PN series relies on the study of open orbits, our hypergeometric representations provide an algorithm which is very efficient when implemented in a Mathematica code to describe PN expansions of bounded orbits. Indeed, here we compute the gravitational waveform and the energy flux, in the case of circular orbits, up to the 30th PN order [^2] The ancillary mathematica files coming with this paper should help the reader to check our computations.. This result will be tested against the 22nd PN order computation available in literature Fujita:2012cm , and results based on MST and numerical methods from the Black Hole Perturbation Toolkit (BHTK) BHPToolkit . The analytic expressions of all ingredients contributing to the 30PN expansion of the wave form and energy flux are collected in ancillary mathematical files. Numerical studies of unbounded systems are postponed to future investigations.
This is the plan of the paper: in Section 2 we derive the PM and PN expansion of the confluent Heun function. In Section 3 we apply the results of Section 2 to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section 4 we test our results against those obtained via the MST method and numerical integration. In Section 5 we extend the results to include the case of the non-confluent Heun function. Section 6 contains our conclusions.
2 The confluent Heun equation
The confluent Heun differential equation
$g^{\prime\prime}(t)+\left(\frac{\beta_{0}}{t}+\frac{\beta_{1}}{t-1}+\lambda\right)g^{\prime}(t)+\frac{\alpha_{1}t-\alpha_{0}}{t(t-1)}g(t)=0$ (1)
is an ordinary differential equation of second order with two regular singularities at $t=0,1$ and an irregular one at infinity. A basis of solutions is given by
${\rm HeunC}[\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},\lambda,t]$ (2)
$t^{1-\beta_{0}}\text{HeunC}[\alpha_{0}+(1-\beta_{0})(\lambda-\beta_{1}),\alpha_{1}+\lambda(1-\beta_{0}),2-\beta_{0},\beta_{1},\lambda,t]$
with ${\rm HeunC}$ the confluent Heun function.
2.1 Confluent Heun-Gauge dictionary
The confluent Heun equation can be viewed as the quantum version of the Seiberg-Witten (SW) curve governing the dynamics of a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with three massive flavours ($m_{i}$)transforming in the fundamental representation of the gauge group (with gauge coupling $x$) Bianchi:2021mft ; Bianchi:2021xpr living on a curved Nekrasov-Shatashvilli background Nekrasov:2009rc . This suggests a a change of variables in the Heun equation (1) in terms of the parameters of the SW quantum curve and its Coulomb branch parameter $u$
$\alpha_{0}$ $=$ $u-{\textstyle\frac{1}{4}}+m_{1}-m_{1}^{2}+(m_{1}+m_{3}-1)x\qquad,\qquad\alpha_{1}=\left(m_{1}+m_{3}-1\right)x$
$\beta_{0}$ $=$ $1-m_{1}+m_{2}\qquad,\qquad\beta_{1}=1-m_{1}-m_{2}\qquad,\qquad\lambda=-x$ (3)
which will simplify the computations which we are going to perform later in the paper. Defining
$G_{1}(z)=z^{-{\alpha_{1}\over\lambda}}g({\textstyle\frac{1}{z}})$ (4)
the Heun equation (1) can be put into the form [^3] Alternatively, equation (8) can be written in the normal Schrödinger like form $\Psi^{\prime\prime}(z)+Q_{\rm gauge}(z)\Psi=0\qquad,\qquad G_{1}(z)=e^{{\frac{x}{2z}}}z^{m_{3}-1}(1-z)^{m_{1}+m_{2}-1\over 2}\Psi(z)$ (5) with $Q_{\rm gauge}={-}\frac{x^{2}}{4z^{4}}{+}\frac{xm_{3}}{z^{3}}{+}\frac{1{-}\left(m_{1}{-}m_{2}\right)^{2}}{4(z{-}1)z}+\frac{1{-}\left(m_{1}{+}m_{2}\right){}^{2}}{4(z-1)^{2}z}+\frac{u{-}{\textstyle\frac{1}{4}}{+}{\textstyle\frac{x}{2}}\left(m_{1}{+}m_{2}{-}1\right)}{(z{-}1)z^{2}}$ (6) Notice that also the conformal Ward identity of Liouville theory, the so called BPZ equation, is a confluent Heun equation. The normal form of such equation matches again with (5) and (6)after the identifications $m_{1}=p_{0}-k_{0}\qquad,\qquad m_{2}=-p_{0}-k_{0}\qquad,\qquad m_{3}=c\qquad,\qquad x=-x_{\rm CFT}$ (7) with $k_{0}$, $p_{0}$, $c$ and $x_{\rm CFT}$ the CFT variables in Consoli:2022eey ; Fucito:2023afe .
$G_{1}^{\prime\prime}(z)+G_{1}^{\prime}(z)\left(\frac{1{-}m_{1}{-}m_{2}}{z{-}1}{-}\frac{2\left(m_{3}{-}1\right)}{z}{+}\frac{x}{z^{2}}\right)$ (8)
$\qquad{+}G_{1}(z)\left(\frac{{\textstyle\frac{1}{4}}{-}u{+}m_{3}(m_{3}{-}1)}{z^{2}}{+}\frac{u+\frac{3}{4}-m_{1}-m_{2}-m_{3}+m_{1}m_{2}+m_{1}m_{3}+m_{2}m_{3}}{(z{-}1)z}\right)=0$
where now the irregular point is located at the origin. Alternatively, one writes
$G_{0}(z)=G_{1}({\textstyle\frac{x}{z}})$ (9)
and setting $y=x/z$ the confluent Heun equation (8) becomes
$G_{0}^{\prime\prime}(y){+}G_{0}^{\prime}(y)\left(\frac{1{-}m_{1}{-}m_{2}}{y{-}x}{+}\frac{m_{1}{+}m_{2}{+}2m_{3}{-}1}{y}{-}1\right)$ (10)
$\qquad\qquad{+}G_{0}(y)\left(\frac{u{+}{\textstyle\frac{3}{4}}{+}\delta}{(x-y)y}{+}\frac{1{+}\delta{+}m_{3}(m_{3}{-}1)}{y^{2}}\right)=0$
where $\delta=m_{1}m_{2}+m_{1}m_{3}+m_{2}m_{3}-m_{1}-m_{2}-m_{3}$. We will use both descriptions of the confluent Heun equation, and we will refer to them as CHE. In particular, we will consider the expansions of the two equations in the limit where $x$ is taken small but keeping finite either $z$ or $y=x/z$. The two expansions will correspond in the gravity context to the Post-Minkowskian and Post-Newtonian approximations, so we will refer to $G_{0}$ and $G_{1}$ as the PM and PN representations of the confluent Heun function respectively. In particular, setting $x\to 0$, (10) and (8) reduce to hypergeometric equations with solutions
${\rm PM}:$ $\qquad G^{0}_{\alpha}(y)\underset{x\to 0\atop y~{}{\rm finite}}{\approx}H^{0}_{\alpha}(y)$
${\rm PN}:$ $\qquad G^{1}_{\alpha}(z)\underset{x\to 0\atop z~{}{\rm finite}}{\approx}H^{1}_{\alpha}(z)$ (11)
with $\alpha=\pm$ labeling the two independent solutions,
$H^{0}_{\alpha}(y)$ $=$ $y^{\frac{1}{2}-\alpha a-m_{3}}\,_{1}F_{1}\left({\textstyle\frac{1}{2}}-\alpha a-m_{3};1-2\alpha a;y\right)$
$H^{1}_{\alpha}(z)$ $=$ $z^{-\frac{1}{2}+\alpha a+m_{3}}\,_{2}F_{1}\left(\frac{1}{2}+\alpha a-m_{1},\frac{1}{2}+\alpha a-m_{2};1+2\alpha a;z\right)$ (12)
and $a\underset{x\to 0}{\approx}\sqrt{u}$. On the other hand, at order $k$ in $x$, we will look for solutions of the form
${\rm PM}_{k}:$ $\qquad G^{0}_{\alpha}(y)=y^{\frac{1}{2}-\alpha a-m_{3}}\sum_{i=0}^{k}\sum_{j=0}^{\infty}c^{PM}_{ij}(\alpha a)x^{i}y^{j-i}$
${\rm PN}_{k}:$ $\qquad G^{1}_{\alpha}(z)=z^{-\frac{1}{2}+\alpha a+m_{3}}\sum_{j=0}^{k}\sum_{i=0}^{\infty}c^{PN}_{ij}(\alpha a)z^{i-j}x^{j}$ (13)
where $a(u)$ is a function of $u$. Writing
$u=a^{2}+\sum_{i=1}^{k}u_{i}(a)x^{i}$ (14)
one can determine the coefficients $c^{PM}_{ij}(\alpha a)$, $c^{PN}_{ij}(\alpha a)$ and $u_{i}(a)$ by solving the CHE equations order by order in $x$. We notice that $a$ codify the monodromy of the solutions around the irregular point $z=0$. Moreover, since $G^{1}_{\alpha}(z)$ and $G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ are solutions of the same equation and have the same monodromy around $z=0$ they should be proportional to each other. We conclude therefore that in the overlap region, the so called near zone, where $x\ll z\ll 1$
${\rm Near~{}zone}:\qquad G^{1}_{\alpha}(z)=g_{\alpha}(x)G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ (15)
The main result of this Section is that, the expansions PMk and PNk in (2.1) can be exactly resumed and written in terms of a single hypergeometric function. This will provide us with a first principle proof of the connection formulae for the confluent Heun functions, where the whole non-triviality of the Heun equation is codified into a single function $a(u)$.
For concreteness, we will focus on the domain $z\in[0,1]$. The general solution $G_{1}(z)=\sum_{\alpha}c_{\alpha}G^{1}_{\alpha}(z)$ behaves near the boundaries as
$G_{1}(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{2m_{3}}e^{x\over z}(1+\ldots)$
$G_{1}(z)$ $\underset{z\to 1}{\approx}$ $D_{-}(1+\ldots)+D_{+}(1-z)^{m_{1}+m_{2}}(1+\ldots)\,\,\,\,\,\,\,\,\,$ (16)
where the dots stand for subleading terms in these limits. Specifying boundary conditions corresponds to fix two linear combinations made out of the four coefficients $B_{\pm}$, $D_{\pm}$. We will derive connection formulae relating the coefficients $B_{\pm}$ to $D_{\pm}$ and viceversa.
2.2 PM approximation: exterior region
In this subsection we consider the limit $x\ll 1$ keeping $y$ finite. This corresponds to the PM approximation and we refer to this region as the exterior region: the reason of it will be clear in the next Section from the dictionary with the gravity side (see (45)).
We denote the basic solution as $G_{\alpha}^{0}({\textstyle\frac{x}{z}})$ and at leading order in $x$ we take
$G^{0}_{\alpha}(y)=H^{0}_{\alpha}(y)$ (17)
For higher orders in $x$, proceeding by trial and error, we come to the following ansätze
$G^{0}_{\alpha}(y)=P_{0}(y)\,H^{0}_{\alpha}(y){+}\widehat{P}_{0}(y)\,yH^{0}_{\alpha}{}^{\prime}(y)$ (18)
where the prime stands for derivation and
$P_{0}(y)$ $=$ $1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}$
$\widehat{P}_{0}(y)$ $=$ $\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}$ (19)
are polynomials of order $k$ in $x$ [^4] For $k=1$ M.Billò, M.L.Frau and A.Lerda first obtained (18) resumming the contributions of the Young tableaux entering the partition function of the supersymmetric gauge theory.. We plug this ansätze into (8) and use the differential equation to express higher derivatives of $H^{0}_{\alpha}$ in terms of $H^{0}_{\alpha}$ and its first derivative themselves. Therefore one finds a system of algebraic linear equations that determines the coefficients $c_{ij}$, $\widehat{c}_{ij}$ and $u_{i}$ (whose complete forms are given in (76)). Setting $c_{00}=1$, one finds, for the first term in the recursion
$c_{10}$ $=$ $u_{1}=\frac{1}{2}(1-m_{1}-m_{2}-m_{3})-{2m_{1}m_{2}m_{3}\over 4a^{2}-1}$
$\widehat{c}_{10}$ $=$ $\partial_{m_{3}}u_{1}=-\frac{1}{2}-{2m_{1}m_{2}\over 4a^{2}-1}$ (20)
It is easy to check that $c_{ij}$, $\widehat{c}_{ij}$ are invariant under $a\to-a$. This is true at any PM order, so all the dependence on $\alpha$ of the solution is codified in $H^{0}_{\alpha}$. The results (20) are special cases of more general relations that determine the coefficients $c_{i,i-1}$, $\widehat{c}_{i,i-1}$ in terms of $u$
$c_{i,i-1}=u_{i}\,,$
$\gamma_{m}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}={\rm exp}\left\{\sum_{i=1}^{\infty}{\partial_{m_{3}}u_{i}\over i}x^{i}\right\}\quad$ (21)
The combination in the exponential will be denoted in the following as
${\cal F}_{\rm inst}(a,x)\equiv-\sum_{i=1}^{\infty}{u_{i}(a)\over i}x^{i}$ (22)
The subscript "inst" reminds the reader that this quantity is interpreted as the instanton prepotential in the gauge theory description of the Heun equation. Indeed in the limit $z\to 0$, where one of the two gauge couplings is turned off, the quiver theory reduces to a ${\cal N}=2$ supersymmetric SU(2) gauge theory with three massive fundamental flavours with prepotential ${\cal F}_{\rm inst}$.
We notice that (18) is exact in $y$ at the k-th PM order, so we can extrapolate it to the region of large $y$ using the standard hypergeometric transformation rule
$H^{0}_{\alpha}(y)=\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{H}^{0}_{\alpha^{\prime}}(y)$ (23)
with
$\widetilde{H}^{0}_{\alpha}(y)=y^{-m_{3}(1+\alpha)}e^{y(1+\alpha)\over 2}{}_{2}F_{0}\left(\frac{1}{2}-a+\alpha m_{3},\frac{1}{2}+a+\alpha m_{3},\frac{\alpha}{y}\right)$ (24)
and
$B_{\alpha\alpha^{\prime}}=e^{\frac{{\rm i}\pi}{2}\left(1-\alpha^{\prime}\right)\left({1\over 2}-\alpha a-m_{3}\right)}\frac{\Gamma(1-2\alpha a)}{\Gamma\left(\frac{1}{2}-\alpha a-\alpha^{\prime}m_{3}\right)}$ (25)
We can therefore introduce an equivalent basis of solutions, naturally defined in the far region.
$\widetilde{G}^{0}_{\alpha}(y)=P_{0}(y)\widetilde{H}^{0}_{\alpha}(y){+}\widehat{P}_{0}(y)y\widetilde{H}^{0}_{\alpha}{}^{\prime}(y)$ (26)
related to the previous one by the connection formulae
$G^{0}_{\alpha}(y)=\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}(y)$ (27)
2.3 PN approximation: interior region
In this subsection we consider the limit $x\ll 1$ keeping $z$ finite. This corresponds to the PN approximation and we refer to this region as the interior region. The analysis proceeds mutatis mutandis following the same steps of the previous subsection, exchanging the roles of $y$ and $z$.
We denote by $G^{1}_{\alpha}(z)$ the basic solutions and at leading order in $x$ we take
$G^{1}_{\alpha}(z)=H^{1}_{\alpha}(z)$ (28)
Higher orders follow from the ansätze
$G^{1}_{\alpha}(z)=P_{1}(z)\,H^{1}_{\alpha}(z){+}\widehat{P}_{1}(z)\,zH^{1}_{\alpha}{}^{\prime}(z)$ (29)
where the prime stands for derivation and
$P_{1}(z)$ $=$ $1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}$
$\widehat{P}_{1}(z)$ $=$ $\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}$ (30)
Setting $d_{00}=1$, one finds that the first non-trivial coefficients are
$d_{01}=\frac{1}{2}+\frac{2m_{3}\left(1-m_{3}\right)}{4a^{2}-1},\quad\widehat{d}_{01}=-\widehat{d}_{11}=\frac{2m_{3}}{4a^{2}-1}$ (31)
We notice that the results are invariant under $a\to-a$. This is true at any PN order, so all the dependence on $\alpha$ of the solution is codified in $H^{1}_{\alpha}$.
Proceeding as before, we can extrapolate the PN representation to the region $z\approx 1$ using the hypergeometric transformation rule
$H^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{H}^{1}_{\alpha^{\prime}}(z)$ (32)
with
$\widetilde{H}^{1}_{\alpha}(z)=z^{a{+}m_{3}{-}\frac{1}{2}}(1{-}z)^{\frac{1}{2}\left(m_{1}{+}m_{2}\right)\left(1{+}\alpha\right)}\,_{2}F_{1}\left(\frac{1}{2}{+}a{+}\alpha m_{1},\frac{1}{2}{+}a{+}\alpha m_{2};1{+}\alpha\left(m_{1}{+}m_{2}\right);1{-}z\right)$ (33)
and
$F_{\alpha\alpha^{\prime}}=\frac{\Gamma(1{+}2\alpha a)\Gamma\left(-\alpha^{\prime}\left(m_{1}+m_{2}\right)\right)}{\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{1}\right)\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{2}\right)}$ (34)
An equivalent basis of solutions defined in the near horizon region is therefore given by
$\widetilde{G}^{1}_{\alpha}(z)=P_{1}(z)\,\widetilde{H}^{1}_{\alpha}(z){+}\widehat{P}_{1}(z)\,z\,\widetilde{H}^{1}_{\alpha}{}^{\prime}(z)$ (35)
satisfying the connection formulae
$G^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}(z)$ (36)
2.4 Boundary conditions and connection formulae
The functions $G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ and $G^{1}_{\alpha}(z)$ are solutions of the same differential equation and have the same monodromy around $z=0$, so they should be proportional to each other. We denote by $g_{\alpha}(x)$ their ratio
$g_{\alpha}(x)={G^{1}_{\alpha}(z)\over G^{0}_{\alpha}({\textstyle\frac{x}{z}})}=x^{\alpha a{+}m_{3}{-}\frac{1}{2}}\left[1{+}x\left({\textstyle\frac{1}{4}}{-}{\textstyle\frac{m_{3}\left(m_{1}{+}m_{2}{+}2m_{3}{-}2\right){-}m_{1}m_{2}}{4a^{2}-1}}{+}{\textstyle\frac{8\,\alpha\,a\,m_{1}m_{2}m_{3}}{\left(4a^{2}{-}1\right)^{2}}}\right)+\ldots\right]$ (37)
Expanding for small $z$ and $y$ one can check explicitly that this ratio depends only on $x=zy$. Furthermore one finds
$\gamma(x)={g_{+}(x)\over g_{-}(x)}=x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}(a,x)}$ (38)
with ${\cal F}_{\rm inst}(a,x)$ defined in (22). Collecting all the different pieces together, we have at our disposal four different bases of solutions, so that the general solution can be written as
$G_{1}(z)$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)G^{0}_{\alpha}({\textstyle\frac{x}{z}})+\ldots=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}({\textstyle\frac{x}{z}})+\ldots$ (39)
$=$ $\sum_{\alpha}c_{\alpha}G^{1}_{\alpha}(z)+\ldots=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}(z)+\ldots$
The first line corresponds to the PM approximation where $x$ is taken small keeping $y$ finite. The second equality in this line is obtained using the connection formulae (27). Similarly, the second line corresponds to the PN approximation where $x$ is taken small keeping $z$ finite and the right hand side follows from (36). Finally dots stand for terms of order $x^{k+1}$ in the corresponding limits. Expanding at the leading order the two right hand sides in (39) and using (21), one finds the asymptotics
$G_{1}(z)$ $\underset{z\to 0}{\approx}$ $\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}({\textstyle\frac{x}{z}})^{-m_{3}(1+\alpha^{\prime})}e^{x(1+\alpha^{\prime})\over 2z}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$
$G_{1}(z)$ $\underset{z\to 1}{\approx}$ $\sum_{\alpha,\alpha^{\prime}}c_{\alpha}F_{\alpha\alpha^{\prime}}(1{-}z)^{{1{+}\alpha^{\prime}\over 2}(m_{1}{+}m_{2})}\,h_{\alpha^{\prime}}$ (40)
with
$h_{\alpha^{\prime}}=P_{1}(1)\,{+}{1{+}\alpha^{\prime}\over 2}(m_{1}{+}m_{2})\widehat{P}_{1}{}^{\prime}(1)\,$ (41)
Once again the ratio of these coefficients takes a simple form
$\frac{h_{+}}{h_{-}}=e^{\partial_{m_{1}+m_{2}}{\cal F}_{\rm inst}}$ (42)
Comparing against (16) one finds the asymptotic coefficients
$B_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}x^{-m_{3}(1+\alpha^{\prime})}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$
$D_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c_{\alpha}F_{\alpha\alpha^{\prime}}h_{\alpha^{\prime}}$ (43)
3 Black hole perturbation
In this Section we apply the results obtained in the previous one to the study of gravitational waves in the Schwarzschild geometry.
3.1 Heun gravity correspondence
BHs perturbations are typically described by equations of the confluent Heun type. For example, the Teukolsky equation for spin $s$ perturbations in the Schwarzschild geometry is given by
${1\over\Delta(r)^{s}}{d\over dr}\left[\Delta(r)^{s+1}R^{\prime}(r)\right]{+}\left(\frac{\omega^{2}r^{4}{-}2{\rm i}s(r{-}M)\omega r^{2}}{\Delta(r)}{+}4{\rm i}s\omega r{-}(\ell{-}s)(\ell{+}s{+}1)\right)R(r)=0$ (44)
with $\Delta=r(r-2M)$. We are interested in the $\psi_{4}$ fundamental mode Newman:1961qr ; Teukolsky:1973ha corresponding to $s=-2$. (44) can be put into the CHE form (8) after the following identifications
$R(z)$ $=$ $e^{-{x\over 2z}}(1-z)^{1-{m_{1}+m_{2}\over 2}}({\textstyle\frac{x}{z}})^{1+m_{3}}G_{1}(z)\qquad,\quad z={2M\over r}\qquad,\qquad x=4iM\omega$ (45)
$m_{1}$ $=$ $2{\rm i}M\omega\quad,\quad m_{2}=m_{3}^{*}=-2+2{\rm i}M\omega\quad,\quad u=(\ell+{\textstyle\frac{1}{2}})^{2}+10{\rm i}M\omega-4\omega^{2}M^{2}$
We denote by $R_{\rm in}(z)$ a solution satisfying incoming boundary conditions at the horizon and by $R_{\rm up}(z)$ that one satisfying upgoing boundary conditions at infinity, i.e.
$R_{\rm in}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm in}_{-}e^{-{x\over 2z}}z^{-1-m_{3}}+B^{\rm in}_{+}e^{{x\over 2z}}z^{m_{3}-1}&{}{}{}{}{}{}{}&\quad z\to 0\\ D^{\rm in}_{-}(1-z)^{1-{m_{1}+m_{2}\over 2}}&&\quad z\to 1\\ \end{array}\right.$ (48)
$R_{\rm up}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm up}_{+}e^{{x\over 2z}}z^{m_{3}-1}&{}{}{}{}{}{}{}&\quad z\to 0\\ D^{\rm up}_{-}(1-z)^{1-{m_{1}+m_{2}\over 2}}+D^{\rm up}_{+}(1-z)^{1+{m_{1}+m_{2}\over 2}}&&\quad z\to 1\\ \end{array}\right.$ (51)
where we have set to zero the coefficients $D^{\rm in}_{+}$ and $B^{\rm up}_{-}$. In radial coordinates $R_{\rm in,up}(r)\equiv R_{\rm in,up}\left[z(r)\right]$. In terms of these solutions the Green function reads
$G(r,r^{\prime})={1\over W}\left\{\begin{array}[]{ccc}R_{\rm in}(r^{\prime})R_{\rm up}(r)&{}{}{}{}{}{}{}\hfil&r^{\prime}<r\\ R_{\rm in}(r)R_{\rm up}(r^{\prime})&{}{}{}{}{}{}{}\hfil&r<r^{\prime}\\ \end{array}\right.$ (52)
with
$W={R_{\rm in}(r)R^{\prime}_{\rm up}(r)-R^{\prime}_{\rm in}(r)R_{\rm up}(r)\over\Delta(r)}={{\rm i}\omega\over 2M^{2}}B^{\rm in}_{-}B^{\rm up}_{+}$ (53)
a constant built out of the Wronskian which is conveniently computed for $r\to\infty$, since its first derivative is zero as it follows using (44). In the following we use as independent variables the two dimensionless combinations
$y$ $=$ $2{\rm i}\omega r\qquad,\qquad x=4{\rm i}\omega M$ (54)
The incoming and upgoing solutions can be written as linear combinations
$R_{\rm in,up}(y)=\sum_{\alpha}c^{\rm in,up}_{\alpha}g_{\alpha}R_{\alpha}(y)$ (55)
of the fundamental solutions $R_{\alpha}(y)$. Since we are mainly interested in the physics at large distances $z\ll 1$, it is convenient to work with the PM representation and write the two independent solutions of the Teukolsky equation as
$R_{\alpha}(y)=e^{-{y\over 2}}\left(1-\frac{x}{y}\right)^{2{-}{x\over 2}}y^{{-}1{-}{x\over 2}}G^{0}_{\alpha}(y)$ (56)
$c^{\rm in,up}_{\alpha}$ are some coefficients, chosen such that the incoming boundary conditions at the horizon or upgoing at infinity are satisfied, i.e.
$D^{\rm in}_{+}=B^{\rm up}_{-}=0$ (57)
with asymptotic coefficients given by (16), (45)
$B^{\rm in,up}_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c^{\rm in,up}_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}x^{1-\alpha^{\prime}m_{3}}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$
$D^{\rm in,up}_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c^{\rm in,up}_{\alpha}F_{\alpha\alpha^{\prime}}h_{\alpha^{\prime}}e^{-{x\over 2}}\,x^{1+m_{3}}$ (58)
We are interested in the PN expansion of the solutions, since we want to describe binary systems in the limit where the distance between the two objects is large and velocities small. For example, for circular orbits with tangential velocity $v$, this corresponds to the limit where $v$ is small, $y\sim v$ and $x\sim v^{3}$. Therefore the PN expansion can be written as
$R_{\alpha}(y)=y^{{3\over 2}-\alpha a}\sum_{n=0}^{\infty}R_{n}(\alpha a,y)$ (59)
with
$R_{0}(a,y)=1$
$R_{1}(a,y)=\frac{2y}{1-2a}$
$R_{2}(a,y)=\frac{(17-2a)y^{2}}{16(2a-1)(a-1)}+\frac{(2a-3)x}{4y}$
$R_{3}(a,y)=\frac{(7-2a)y^{3}}{8(2a-1)(2a-3)(a-1)}+\frac{(11-6a)x}{4(2a-1)}$
$R_{4}(a,y)=\frac{\left(4a^{2}{-}72a{+}163\right)y^{4}}{512(a{-}2)(a{-}1)(2a{-}3)(2a{-}1)}$
$+\frac{\left({-}4a^{2}{+}48a{-}119\right)yx}{64(2a{-}1)(a{-}1)}{+}\frac{\left(8a^{3}{-}4a^{2}{-}18a{+}9\right)x^{2}}{64y^{2}(a{+}1)}$ (60)
and
$a(\ell)=\ell+\frac{1}{2}+\frac{\left(15\ell^{4}+30\ell^{3}+28\ell^{2}+13\ell+24\right)x^{2}}{8\ell(\ell+1)(2\ell+1)\left(4\ell^{2}+4\ell-3\right)}+\ldots$ (61)
3.2 The incoming solution
The incoming solution is obtained by requiring no outgoing waves at the horizon, i.e. by setting $D^{\rm in}_{+}=0$. According to (58), such condition leads to
$\sum_{\alpha=\pm}c^{\rm in}_{\alpha}F_{\alpha+}=0$ (62)
This determines the ratio of the two coefficients to be
$\ {c^{\rm in}_{+}\over c^{\rm in}_{-}}=-{F_{-+}\over F_{++}}=\frac{\Gamma({-}2a)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{1}{+}a\right)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{2}{+}a\right)}{\Gamma(2a)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{1}{-}a\right)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{2}{-}a\right)}$ (63)
On the other hand, using the first of (40), one finds the asymptotic behaviour at infinity
$R_{\rm in}(y)\underset{y\to\infty}{\approx}\sum_{\alpha,\alpha^{\prime}}c^{\rm in}_{\alpha}g_{\alpha}B_{\alpha\alpha^{\prime}}e^{\alpha^{\prime}y\over 2}y^{1-\alpha^{\prime}m_{3}}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$ (64)
Comparing against the first line of $R_{\text{in}}$ in (51), one finds
$B^{\rm in}_{\alpha^{\prime}}=\sum_{\alpha}c^{\rm in}_{\alpha}g_{\alpha}B_{\alpha\alpha^{\prime}}x^{1-\alpha^{\prime}m_{3}}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$ (65)
We notice that the incoming boundary condition determines only the ratio $c^{\rm in}_{+}/c^{\rm in}_{-}$, leaving undetermined the overall normalization $c^{\rm in}_{-}$. This dependence however cancels out in the Green function after dividing by the factor $W$ (53). It is convenient therefore to define a ratio which does not depend on overall normalizations
$\mathfrak{R}_{\rm in}(y)={R_{\rm in}(y)\over xB^{\rm in}_{-}}=C_{\rm in}\left[R_{-}(y)-\gamma{\textstyle\frac{F_{-+}}{F_{++}}}R_{+}(y)\right]$ (66)
with
$C_{\rm in}={x^{x\over 2}\over B_{--}}\left(1-\gamma{\textstyle\frac{F_{-+}B_{+-}}{F_{++}B_{--}}}\right)^{-1}\underset{x\to 0}{\approx}(-1)^{\ell+1}\frac{\Gamma\left(\ell{-}1\right)}{\Gamma(2\ell{+}2)}+\ldots$ (67)
We notice that in the PN limit $R_{\alpha}(y)\sim y^{{3\over 2}-\alpha a}$ and $\gamma\sim x^{2a}$, so the $\gamma R_{+}$-contribution in (66) is suppressed by an extra factor $(x/y)^{2a}$ with respect to $R_{-}$. The incoming solution is therefore dominated by the $R_{-}$ component. Using (61) one finds the PN expansion
$R_{-}(y)=y^{a+{3\over 2}}\left[1{+}\frac{y}{1{+}\ell}{+}\left(\frac{(9{+}\ell)y^{2}}{8(1{+}\ell)(3{+}2\ell)}{-}\frac{(\ell{+}2)x}{2y}\right){+}\left(\frac{(\ell{+}4)y^{3}}{8(\ell{+}1)(\ell{+}2)(2\ell{+}3)}{-}\frac{(7{+}3\ell)x}{4(\ell+1)}\right)\right.$
$\left.+\left(\frac{\left(50{+}19\ell{+}\ell^{2}\right)y^{4}}{128(\ell{+}1)(\ell{+}2)(2\ell{+}3)(2\ell{+}5)}{+}\frac{(\ell^{2}-1)(\ell+2)x^{2}}{4(2\ell-1)y^{2}}-\frac{(\ell+4)(\ell+9)xy}{16(\ell+1)(2\ell+3)}\right)+\ldots\right]$ (68)
It is important to observe that at higher orders both $R_{\pm}(y)$ exhibit poles in the limit where $\ell\to\mathbb{Z}$, but one can check that they cancel against each other in the combination (66). For example, for $\ell=2$ the poles in the two components first appear at order $v^{16}$.
3.3 The upgoing solution
Upgoing boundary conditions correspond to setting $B^{\rm up}_{-}=0$, that according to (58) leads to
${c_{-}^{\rm up}\over c_{+}^{\rm up}}=-\gamma{B_{+-}\over B_{--}}$ (69)
Following the same passages of the incoming solution, we can write the asymptotic behaviour of $R_{\text{up}}(y)$ at infinity, which is the same as (64) but with $c^{\text{up}}_{\alpha}$ in place of $c^{\text{in}}_{\alpha}$. This implies that also the coefficients $B_{\alpha}^{\text{up}}$ take the same form of (65) with the same substitution $c^{\text{in}}_{\alpha}\to c^{\text{up}}_{\alpha}$. Again we introduce the ratio
$\mathfrak{R}_{\rm up}(y)=(2M)^{3}{R_{\rm up}(y)\over B^{\rm up}_{+}}=C_{\rm up}\left[R_{+}(y)-{\textstyle\frac{B_{+-}}{B_{--}}}R_{-}(y)\right]$ (70)
with
$C_{\rm up}$ $=$ ${(2M)^{3}e^{\partial_{m_{3}}{\cal F}_{\rm inst}}\over x^{3+\frac{x}{2}}B_{++}}\left(1-{\textstyle\frac{B_{-+}B_{+-}}{B_{++}B_{--}}}\right)^{-1}=\frac{(2M)^{3}}{x^{3+\frac{x}{2}}}\frac{\Gamma(2-\ell)}{\Gamma(-2\ell)}+\ldots$ (71)
We notice that the leading term on the right hand side is finite in the limit where $\ell$ becomes an integer with $\ell\geq 2$. The same result is obtained by setting $\ell\to\mathbb{N}$ from the very beginning and sending then $x\to 0$. In the PN limit, the solution is dominated by the $R_{+}(y)$ component with PN expansion
$R_{+}(y)=y^{{3\over 2}-a}\left[1{-}\frac{y}{\ell}{+}\left(\frac{(\ell-1)x}{2y}{-}\frac{(\ell{-}8)y^{2}}{8l(2\ell{-}1)}\right){+}\left(-\frac{(3\ell{-}4)x}{4\ell}{+}\frac{(\ell{-}3)y^{3}}{8(\ell{-}1)\ell(2\ell{-}1)}\right)\right.$
$\left.+\left(\frac{(\ell{-}1)\ell(\ell{+}2)x^{2}}{4(2\ell{+}3)y^{2}}-\frac{(\ell{-}8)(\ell{-}3)xy}{16\ell(2\ell{-}1)}+\frac{\left(32{-}17\ell{+}\ell^{2}\right)y^{4}}{128(\ell{-}1)\ell(2\ell{-}3)(2\ell{-}1)}\right)\right]$ (72)
It is important to observe that again the poles in $R_{\pm}(y)$ in the limit where $\ell\to\mathbb{Z}$ cancel against each other in the combination (70). For $\ell=2$ the poles in the two components first appear at order $v^{7}$.
4 Wave form and energy flux at 30PN
In this Section we compute the wave form and the energy flux produced by a circular EMRI binary system at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the light particle. The results will be compared against the 22PN results available in the literature Fujita:2012cm and the numerical computations using the package Teukolsky of the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST method Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk and on a numerical integrations of the differential equation. We stress the fact that by 30PN, we mean that the polynomials coefficients $P_{0}$, $\widehat{P}_{0}$ and the exponent $a$, defining the solution are truncated at order 30PN, but the hypergeometric functions and the $c^{\rm in}_{\alpha}(a)$ coefficients specifying the solution are kept exact, so an infinite number of terms beyond 30PN are taken into account. In particular, the infinite tower of tail contributions (log divergences, $\pi$’s and transcendental numbers) are resummed into exponentials and Gamma functions. The observables we consider are the $(\ell,m)$ harmonic modes of the incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system.
Since for circular orbits $v^{2}=M/r$, and $r$ should be taken bigger than the innermost stable circular orbit (ISCO) radius $r_{\text{ISCO}}=6M$, the domain of variability of the tangential velocity is $v\in[0,{\textstyle\frac{1}{\sqrt{6}}}]$. We find that the first 12 digits of 30PN results are stable for any choice $v$ in this domain, so we will always display this number of digits. The numerical computations and those with the MST method are performed setting in the BHTK Mathematica files: AccuracyGoal = 1000, PrecisionGoal = 70, WorkingPrecision = 80, SetPrecision=80. With this choice of parameters both numerical and MST results agree to a 12 digits precision. In our tables these results will be denoted as BHTK.
4.1 The incoming and upgoing solutions
In this subsection we collect all the ingredients needed in the computations of the incoming, upgoing waveforms at order k=30 in the hypergeometric PM representation. We write555 Here we rewrite the second terms in the right hand sides of (66), (70), as the images under $a\to-a$ of the first terms, using the transformation rules $R_{-}\to R_{+}$, $\gamma\to\gamma^{-1}$, $F_{-\alpha^{\prime}}\leftrightarrow F_{+\alpha^{\prime}}$ and $B_{-\alpha^{\prime}}\leftrightarrow B_{+\alpha^{\prime}}$.
$\mathfrak{R}_{\rm in}(y)$ $=$ $C_{\rm in}R_{-}(y)+(a\to-a)$
$\mathfrak{R}_{\rm up}(y)$ $=$ $C_{\rm up}R_{+}(y)+(a\to-a)$ (73)
with [^6] We notice that $\gamma_{\rm tot}=\gamma{\textstyle\frac{F_{-+}B_{+-}}{F_{++}B_{--}}}=e^{-2\pi{\rm i}a_{D}}=e^{-\partial_{a}({\cal F}_{\rm tree}+{\cal F}_{\rm one-loop}+{\cal F}_{\rm inst})}$ collects the tree level, one-loop and instanton contributions to the dual SW period $a_{D}$.
$R_{\alpha}(y)$ $=$ $e^{-{y\over 2}}\left(1-{\textstyle\frac{x}{y}}\right)^{2{-}{x\over 2}}y^{{-}1{-}{x\over 2}}\left[P_{0}(y)\,H^{0}_{\alpha}(y){+}\widehat{P}_{0}(y)\,yH^{0}_{\alpha}{}^{\prime}(y)\right]$
$H^{0}_{\alpha}(y)$ $=$ $y^{\frac{5}{2}+\frac{x}{2}-\alpha a}\,_{1}F_{1}\left({\textstyle\frac{5}{2}}+{\textstyle\frac{x}{2}}-\alpha a;1-2\alpha a;y\right)$
$C_{\rm in}$ $=$ $x^{x\over 2}e^{-\frac{{\rm i}\pi}{2}(5+2a+x)}{\Gamma(-{\textstyle\frac{3}{2}}+a-{\textstyle\frac{x}{2}})\over\Gamma(1+2a)}\left(1-\gamma_{\rm tot}\right)^{-1}$
$C_{\rm up}$ $=$ ${(2M)^{3}\over x^{3+\frac{x}{2}}\,\gamma_{m}}{\Gamma({\textstyle\frac{5}{2}}-a+{\textstyle\frac{x}{2}})\over\Gamma(1-2a)}\left(1-e^{-2\pi{\rm i}a}\,{\textstyle\frac{\cos{\textstyle\frac{\pi}{2}}(x+2a)}{\cos{\textstyle\frac{\pi}{2}}(x-2a)}}\right)^{-1}$
$\gamma_{m}$ $=$ $1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}\quad,\quad\gamma=x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}}$
$\gamma_{\rm tot}$ $=$ $\gamma\,e^{-2\pi{\rm i}a}\,\frac{\Gamma({-}2a)^{2}\Gamma\left({-}\frac{3}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{+}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{+}a{-}\frac{x}{2}\right)}{\Gamma(2a)^{2}\Gamma\left(-\frac{3}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{1}{2}{-}a{-}\frac{x}{2}\right)\Gamma\left(\frac{5}{2}{-}a{-}\frac{x}{2}\right)}$ (74)
The coefficients $c_{ij},\widehat{c}_{ij}$ in the polynomials
$P_{0}(y)=1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\quad,\quad\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i-1}\widehat{c}_{ij}\,x^{i}y^{j-i}\$ (75)
are determined by the recursion relations
$c_{ij}$ $=$ $\left[A_{1}\,c_{i,j-1}+A_{2}\,c_{i-1,j}+A_{3}\,c_{i-1,j-1}+\widehat{A}_{1}\widehat{c}_{i,j-1}+\widehat{A}_{2}\widehat{c}_{i-1,j}+\widehat{A}_{3}\widehat{c}_{i-1,j-1}\right.$
$\left.+\sum_{s=1}^{k}c_{s,s-1}(C_{1}\,c_{i-s,j-s}+\widehat{C}_{1}\,\widehat{c}_{i-s,j-s})\right]\left[8\Delta_{ij}(\Delta_{ij}^{2}-4a^{2})\right]^{-1}$
$\widehat{c}_{ij}$ $=$ $\left[B_{1}\,c_{i,j-1}+B_{2}\,c_{i-1,j}+B_{3}\,c_{i-1,j-1}+\widehat{B}_{1}\widehat{c}_{i,j-1}+\widehat{B}_{2}\widehat{c}_{i-1,j}+\widehat{B}_{3}\widehat{c}_{i-1,j-1}\right.$
$\left.+\sum_{s=1}^{k}c_{s,s-1}(C_{2}\,c_{i-s,j-s}+\widehat{C}_{2}\,\widehat{c}_{i-s,j-s})\right]\left[8\Delta_{ij}(\Delta_{ij}^{2}-4a^{2})\right]^{-1}$
$u_{i}$ $=$ $c_{i,i-1}$ (76)
starting from $c_{0i}=\delta_{i0}$, $\widehat{c}_{0i}=0$, with
$A_{1}$ $=$ $-8\left(\Delta_{ij}+1\right)\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{A}_{1}=4\left(4a^{2}-(x+5)^{2}\right)\Delta_{i,j}$
$A_{3}$ $=$ $8\Delta_{ij}\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{A}_{3}=4\left(4a^{2}-(x+5)^{2}\right)\left(-\Delta_{i,j}+x-2\right)$
$A_{2}$ $=$ $2\left(3x^{2}{+}6x{-}12a^{2}{-}13\right)\Delta_{ij}{+}8\Delta_{ij}^{3}{-}8(x{-}1)\Delta_{ij}^{2}{+}8a^{2}(x{-}7){-}2(x^{3}{+}3x^{2}{-}13x{-}15),$
$\widehat{A}_{2}$ $=$ $\left(4a^{2}-(x+5)^{2}\right)\left(-2(x-1)\Delta_{ij}+4a^{2}+x^{2}-2x-3\right)$
$B_{1}$ $=$ $-16\left(\Delta_{ij}+1\right)~{},~{}\widehat{B}_{1}=8\Delta_{ij}\left(\Delta_{ij}+x+5\right)~{},~{}\widehat{B}_{3}=8\left(-\Delta_{i,j}^{2}-7\Delta_{ij}+x^{2}+3x-10\right)$
$B_{2}$ $=$ $4\left(-2(x-1)\Delta_{i,j}+4a^{2}+x^{2}-2x-3\right)~{},~{}B_{3}=16\Delta_{ij}$
$\widehat{B}_{2}$ $=$ $2\left(-\Delta_{ij}+x+1\right)\left(-4\Delta_{ij}^{2}-8\Delta_{ij}+12a^{2}+x^{2}+2x-15\right)$
$C_{1}$ $=$ $8\left(\Delta_{ij}-x-5\right)~{},~{}\widehat{C}_{1}=16a^{2}-4(x+5)^{2}~{},~{}C_{2}=16~{},~{}\widehat{C}_{2}=8\left(\Delta_{ij}+x+5\right)$ (77)
and $\Delta_{ij}=i-j$.
Finally one has to compute $a(\ell)$. Given $u_{i}(a)$ by (76), one can invert the series (2.1) to compute $a(u)$ and then set (from (45))
$u=(\ell+{\textstyle\frac{1}{2}})^{2}+{\textstyle\frac{5x}{2}}+{\textstyle\frac{x^{2}}{4}}$ (78)
Inverting the series at high PM order can be time consuming for a Mathematica code. There is an alternative, more direct way to find the function $a(u)$ based on the connection between this quantity and the quantum SW period of a SU(2) theory with three fundamental flavours. The quantum SW period $a$ is known to satisfy the continuous fraction equation Poghosyan:2020zzg
$\frac{xM(a+1)}{P(a+1)-\frac{xM(a+2)}{P(a+2)-\ldots}}+\frac{xM(a)}{P(a-1)-\frac{xM(a-1)}{P(a-2)-\ldots}}-P(a)=0$ (79)
with
$P(a)$ $=$ $a^{2}+ax+2x-\frac{3x^{2}}{4}x-(\ell+{\textstyle\frac{1}{2}})^{2},$
$M(a)$ $=$ $(a-{\textstyle\frac{x}{2}}-{\textstyle\frac{1}{2}})(a-{\textstyle\frac{x}{2}}+{\textstyle\frac{3}{2}})(a+{\textstyle\frac{x}{2}}+{\textstyle\frac{3}{2}})$ (80)
Equation (79) can be solved for $a(\ell)$ order by order in $x$. For example for $\ell=2,3$ one finds
$a(2)$ $=$ ${\textstyle\frac{5}{2}}+{\textstyle\frac{107x^{2}}{840}}-{\textstyle\frac{1695233x^{4}}{148176000}}+{\textstyle\frac{76720109901233x^{6}}{30764716012800000}}-{\textstyle\frac{71638806585865707261481x^{8}}{99644321084605000704000000}}+\ldots$
$a(3)$ $=$ ${\textstyle\frac{7}{2}}+{\textstyle\frac{13x^{2}}{168}}-{\textstyle\frac{10921x^{4}}{4346496}}+{\textstyle\frac{95353832269x^{6}}{493385539046400}}-{\textstyle\frac{23627105510827613x^{8}}{1148838359958761472000}}+\ldots$ (81)
4.2 The luminosity
In this subsection, we collect the formulae needed to evaluate the luminosity at infinity for the case of a particle moving along a circular orbit in a Schwarzschild geometry.
After expanding in harmonics, the corresponding waveform at large distances can be written as
$h=h_{+}-{\rm i}h_{\times}\underset{r\to\infty}{\approx}-{4G\over r}\int{d\omega\over 2\pi}\,\sum_{\ell,m}e^{-{\rm i}\omega(T-R_{*})}\,{Z_{\ell m}(\omega)\over 2({\rm i}\omega)^{2}}Y_{-2}^{\ell m}(\theta,\phi)\;,$ (82)
The harmonic coefficients $Z_{\ell m}$ can be written in terms of derivatives of the incoming solution
$Z_{\ell m}$ $=$ $\mathfrak{R}_{\rm in}(y)\left(\frac{Ex^{2}b_{0\ell m}}{2M^{2}y(x{-}y)}{+}\frac{{\rm i}Jx^{3}\left(y^{2}{+}4y{-}4x\right)b_{1\ell m}}{8M^{3}y^{3}(x{-}y)}{+}\frac{J^{2}x^{4}\left(2x{-}4y{-}y^{2}\right)b_{2\ell m}}{64EM^{4}y^{3}(x{-}y)}\right)$
$+y\partial_{y}\mathfrak{R}_{\rm in}(y)\left(\frac{{\rm i}Jx^{3}b_{1\ell m}}{4M^{3}y^{3}}+\frac{J^{2}x^{4}\left(2x-2y-2y^{2}\right)b_{2\ell m}}{16EM^{4}y^{5}}\right){-}y^{2}\partial_{y}^{2}\mathfrak{R}_{\rm in}(y)\frac{J^{2}x^{4}(x{-}y)b_{2\ell m}}{16EM^{4}y^{5}}$
with
$b^{0}_{\ell m}={\pi\over 2}\sqrt{(\ell-1)\ell(\ell+1)(\ell+2)}\,Y_{0}^{\ell m}({\textstyle\frac{\pi}{2}},0),\,b^{1}_{\ell m}=\pi\sqrt{(\ell-1)(\ell+2)}Y_{-1}^{\ell m}({\textstyle\frac{\pi}{2}},0),\,b^{2}_{\ell m}=\pi\,Y_{-2}^{\ell m}({\textstyle\frac{\pi}{2}},0),$ (84)
and
${}_{s}Y_{\ell m}(\theta,\phi)=e^{im\phi}\sin^{2\ell}\left(\frac{\theta}{2}\right)\sqrt{\frac{(2\ell{+}1)(\ell{-}m)!(\ell{+}m)!}{4\pi(\ell{-}s)!(\ell{+}s)!}}\sum_{r=0}^{\ell{-}s}({-})^{\ell{+}m{-}r{-}s}(^{\ell{-}s}_{s})(^{\ell+s}_{r{+}s{-}m})\cot({\textstyle\frac{\theta}{2}})^{2r{+}s{-}m}$ (85)
the spin-weighted spherical harmonics and
$E={\mu(1-2v^{2})\over\sqrt{1-3v^{2}}}\qquad,\qquad\omega J={\mu mv^{2}\over\sqrt{1-3v^{2}}}$ (86)
the energy and angular momentum of the light particle. Finally, the luminosity, i.e. the amount of energy per unit time emitted by the system towards infinity, is computed by the formula (see Mino:1997bx and references therein)
${d{\cal E}\over dt}=\mu^{2}\sum_{\ell=2}^{\infty}\sum_{m=1}^{\ell}{|Z_{\ell m}|^{2}\over 2\pi\omega^{2}}=\mu^{2}\sum_{\ell=2}^{\infty}\sum_{m=1}^{\ell}\eta_{\ell m}$ (87)
where $\mu$ is the mass of the orbiting body. Rather than the luminosity, we will display its ratio against the luminosity computed in the Newton approximation
$\left({d{\cal E}\over dt}\right)_{N}=\frac{32\mu^{2}v^{10}}{5M^{2}}=\mu^{2}\,\eta_{22}^{0}$ (88)
with $\eta_{22}^{0}$ the contribution of the leading harmonic mode for $v\to 0$.
4.3 Numerical tests
In this subsection we use our formulae computed at order 30PM and set $M=1$. In Figure 1, we plot the real and imaginary parts of the incoming and outgoing solutions obtained with this method for $\ell=2$, $\omega=0.1$.
Figure 1: Plots of incoming and outgoing solutions for $\omega=0.1,\ell=2$. 1a: Re$(\mathfrak{R}_{\text{in}})$, 1b: Im$(\mathfrak{R}_{\text{in}})$, 2a: Re$(\mathfrak{R}_{\text{up}})$, 2b: Im$(\mathfrak{R}_{\text{up}})$.
The results are compared in Table 1 against those obtained by the MST method and numerical integration using the BHTK BHPToolkit . In all the subsequent tables, the displayed digits are those ones that are insensible to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
$\mathfrak{R}_{\text{in}}$ $r=5$ $r=100$
HYP -0.004086908784 -i 0.000826769654 252.949132499007 -i 211.392731187846
BHTK -0.004086908784 -i 0.000826769654 252.949132499009 -i 211.392731187848
$\mathfrak{R}_{\text{up}}$ $r=5$ $r=100$
HYP 2632.048432 -i 840.387207 -15205.329855167 -i 989985.766975648
BHTK 2632.048432 -i 840.387207 -15205.329855168 -i 989985.766975651
Table 1: Comparison between the 30PM representation of confluent Heun solutions and MST/Numerical results for $\omega=0.1,\ell=2$.
In Figure 2 we display the $v$-dependence of the luminosity for the $(\ell,m)=(2,2),(3,3),(2,1)$ harmonic modes computed from (66) at order $k=30$ in the PN approximation.
Figure 2: Luminosity vs velocity. We display the $v$-dependence of the ratio $\eta_{\ell m}/\eta^{0}_{22}$ for the three dominating modes. a: (2,2), b: (3,3), c: (2,1).
The convergence speed of the results is showed in Table 2. The digits in bold style underline the agreement between computations based on the PM hypergeometric representations and MST/numerical results. As expected, the speed of convergence varies with $v$: for $v=0.1$ an agreement on 12 digits is already reached at 10PM order, while for $v=0.4$, contributions at order 30PM are needed to reach a precision on 9 digits.
$k$ $v=0.1$ $v=0.3$ $v=0.4$
5 0.961483706187 0.841473688330 0.887926055076
10 0.961483705502 0.841176615503 0.880582026053
15 0.961483705502 0.841176157786 0.880686334361
20 0.961483705502 0.841176171312 0.880705358744
25 0.961483705502 0.841176171325 0.880705412455
30 0.961483705502 0.841176171325 0.880705411714
BHTK 0.961483705502 0.841176171325 0.880705411627
Table 2: Convergence of the luminosity results. We display the dependence on the PM order $k$ of the ratio $\eta_{2,2}/\eta^{0}_{2,2}$ computed using the hypergeometric representation. The results are compared against MST/ NUM results displayed in the last line.
Finally in Tables 3 and 4 we compare our 30PN results for the energy loss against those obtained by MST/Numerical and by the 22PN computation in Fujita:2012cm .
Method $v=0.1$ $v=0.3$ $v=0.4$
22PN 0.961483705502 0.841176171326 0.880705608981
HYP 0.961483705502 0.841176171325 0.880705411714
BHTK 0.961483705502 0.841176171325 0.880705411627
Table 3: Luminosity $\eta_{2,2}/\eta^{0}_{2,2}$. Comparison between the different methods.
Method $v=0.1$ $v=0.3$ $v=0.4$
22PN 0.974719282291877 0.950003411477562 1.10990619336
HYP 0.974719282291874 0.950003411484251 1.10991197128
BHTK 0.974719282291999 0.950003411484254 1.10991197120
Table 4: Total luminosity $d{\cal E}/dt/(\mu^{2}\eta^{0}_{2,2})$. Comparison between the different methods.
We notice that in the whole range of $v$ the results obtained with the 30PN hypergeometric representations for the total luminosity $d{\cal E}/dt/(\mu^{2}\eta^{0}_{2,2})$ agree with those obtained by MST and numerical integration up to 11-14 digits.
Mathematica codes
Two Mathematica notebooks are given as ancillary files:
  • RinRupdRdt.nb reproduces the results of Tables 1, 3 and 4 at order $k=30$. This code imports the data files P30.m, Pt30.m, gg30.m, gk30.m and a30.m, containing the polynomials $P_{0}(y)$, $\widehat{P}_{0}(y)$, the quantities $\gamma$, $\gamma_{\kappa}$ and the $a$ cycle.
  • PNexpansions.nb computes the PN expansions of $R_{+}(y)$, $\gamma$, $\gamma_{\kappa}$ at any PM order lower than 10. This code imports the data files PNRpp10.m, PNgg10.m, PNgk10.m and PNaa10.m, containing the expansions of the quantities $R_{+}(y)$, $\gamma$, $\gamma_{\kappa}$ and $a$ at order 10PM.
5 The Heun equation
The generalization of the previous results to the case of the non-confluent Heun equation is straightforward. The Heun differential equation is defined as
$G_{1}^{\prime\prime}(z)+\left({\beta_{0}\over z}+{\beta_{1}\over z-1}+{\beta_{2}\over z-\lambda}\right)G_{1}^{\prime}(z)+{z\alpha_{0}\alpha_{1}-\sigma\over z(z-1)(z-\lambda)}G_{1}(z)=0$ (89)
with $\beta_{2}=1+\alpha_{0}+\alpha_{1}-\beta_{0}-\beta_{1}$. The solutions of the Heun equation can be written as linear combinations of
${\rm Heun}_{-}(z)$ $=$ ${\rm Heun}\left[\lambda,\sigma,\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},z\right]$ (90)
${\rm Heun}_{+}(z)$ $=$ $z^{1{-}\beta_{0}}{\rm Heun}\left[\lambda,\sigma{+}\left(1{-}\beta_{0}\right)\left(\beta_{2}{+}\lambda\beta_{1}\right),1{+}\alpha_{0}{-}\beta_{0},1{+}\alpha_{1}{-}\beta_{0},2{-}\beta_{0},\beta_{1},z\right]$
where ${\rm Heun}\left[\lambda,\sigma,\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},z\right]$ is the Heun function which is normalized to one at $z=0$, so that
${\rm Heun}_{-}(z)=1+\ldots\,,\,{\rm Heun}_{+}(z)=z^{1-\beta_{0}}(1+\ldots)$ (91)
The dots stand for corrections in $z$. The expansion coefficients of ${\rm Heun}_{\alpha}(z)$ can be obtained recursively, solving the equation order by order in $z$. The Heun differential equation can be related to the quantum SW curve of a ${\cal N}=2$ supersymmetric gauge theory with four fundamental hypermultiplets. To this aim, we write the six Heun parameters in terms of 4 masses $m_{i}$, a gauge coupling $x$ and a Coulomb branch parameter $u$, defined via the identifications777 Equation (89) can be alternatively written in the normal form (5) with $G_{1}(z)=(1-z)^{-\frac{\beta_{1}}{2}}z^{-\frac{\beta_{0}}{2}}(z-\lambda)^{-\frac{\beta_{2}}{2}}\,\Psi(z)$ (92) and $Q_{\rm gauge}$ $=$ $\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}+m_{4}}{2}}\right){}^{2}}{(z-x)^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{3}-m_{4}}{2}}\right){}^{2}}{z^{2}}+\frac{{\textstyle\frac{1}{4}}-\left({\textstyle\frac{m_{1}+m_{2}}{2}}\right){}^{2}}{(z-1)^{2}}+\frac{m_{3}^{2}+m_{4}^{2}+2m_{1}m_{2}-1}{2(z-1)z}$ (93) $+\frac{(1-x)\left(u+{\textstyle\frac{1}{4}}-\frac{m_{3}^{2}}{2}-\frac{m_{4}^{2}}{2}\right)-{\textstyle\frac{x}{2}}\left(1{-}m_{1}{-}m_{2}\right)\left(1{-}m_{3}{-}m_{4}\right)}{(z-1)z(z-x)}$
$\alpha_{0}$ $=$ $1{-}m_{1}{-}m_{3}\quad,\quad\alpha_{1}=1{-}m_{2}{-}m_{3}\quad,\quad\lambda=x$ (94)
$\beta_{0}$ $=$ $1{-}m_{3}{+}m_{4}\quad,\quad\beta_{1}=1{-}m_{1}{-}m_{2}\quad,\quad\beta_{2}=1-m_{3}-m_{4}$
$\sigma$ $=$ $u(x-1)+\left({\textstyle\frac{1}{2}}-m_{3}\right){}^{2}+x\left({\textstyle\frac{3}{4}}{-}m_{1}{-}m_{2}{-}m_{3}{+}m_{1}m_{2}{+}m_{1}m_{3}{+}m_{2}m_{3}\right)$
The mathematical problem we are interested in is looking for a solution in an interval, let us say $z\in[0,1]$, satisfying some specific boundary conditions. The general solution
$G_{1}(z)=\sum_{\alpha}B_{\alpha}\,{\rm Heun}_{\alpha}(z)$ (95)
behaves, near the boundaries, as
$G_{1}(z)$ $\underset{z\to 1}{\approx}$ $D_{-}(1+\ldots)+D_{+}(1{-}z)^{m_{1}+m_{2}}(1+\ldots)$
$G_{1}(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{m_{3}-m_{4}}(1+\ldots)$ (96)
Specifying boundary conditions corresponds to fix two linear combinations made out of the four coefficients $B_{\pm}$, $D_{\pm}$.
5.1 The hypergeometric representation
As before we start from the PM/PN ansätze
$G^{0}_{\alpha}({\textstyle\frac{x}{z}})g_{\alpha}(x)=G^{1}_{\alpha}(z)+\ldots$ (97)
with
$G^{0}_{\alpha}(y)$ $=$ $P_{0}(y)\,H^{0}_{\alpha}(y){+}\widehat{P}_{0}(y)\,yH^{0}_{\alpha}{}^{\prime}(y)$
$G^{1}_{\alpha}(z)$ $=$ $P_{1}(z)\,H^{1}_{\alpha}(z){+}\widehat{P}_{1}(z)\,zH^{1}_{\alpha}{}^{\prime}(z)$
$H^{0}_{\alpha}(y)$ $=$ $y^{\frac{1}{2}-\alpha a-m_{3}}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}-\alpha a-m_{3},{\textstyle\frac{1}{2}}-\alpha a-m_{4};1-2\alpha a;y\right)$
$H^{1}_{\alpha}(z)$ $=$ $z^{-\frac{1}{2}+\alpha a+m_{3}}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}+\alpha a-m_{1},{\textstyle\frac{1}{2}}+\alpha a-m_{2};1+2\alpha a;z\right)$ (98)
and
$P_{0}(y)$ $=$ $1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\quad,\quad\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i}\widehat{c}_{ij}\,x^{i}y^{j-i}$
$P_{1}(z)$ $=$ $1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}\quad,\quad\widehat{P}_{1}(z)=\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}$
We notice that now only the non-confluent hypergeometric functions are involved. As before, plugging the ansätze (97) into (89) and setting to zero the coefficients of $H^{p}_{\alpha}$ ($p=0,1$) and its derivatives, one finds a linear system of algebraic equations for the coefficients $c_{ij}$, $\widehat{c}_{ij}$, $d_{ij}$, $\widehat{d}_{ij}$. For example, for $k=1$, writing $u=a^{2}+u_{1}x+\ldots$, one finds
$c_{10}$ $=$ $\frac{1}{2}(1-m_{1}-m_{2}-m_{3})-\frac{2m_{1}m_{2}m_{3}}{4a^{2}-1}\quad,\quad\widehat{c}_{11}=-\widehat{c}_{10}=\frac{1}{2}+\frac{2m_{1}m_{2}}{4a^{2}-1}$
$d_{01}$ $=$ $\frac{2\left(m_{3}-1\right)m_{3}m_{4}}{4a^{2}-1}-\frac{m_{4}}{2}\quad,\quad\widehat{d}_{11}=-\widehat{d}_{01}=\frac{2m_{3}m_{4}}{4a^{2}-1}+\frac{1}{2}$
$u_{1}$ $=$ $\frac{2m_{1}m_{2}m_{3}m_{4}}{4a^{2}-1}+\frac{1}{8}\left(4a^{2}-1\right)+\frac{1}{2}\left(1-\sum_{i=1}^{4}m_{i}+\sum_{i<j}^{4}m_{i}m_{j}\right)$ (100)
As in the confluent case, we denote by
$g_{\alpha}(x)={G^{1}_{\alpha}(z)\over G^{0}_{\alpha}({\textstyle\frac{x}{z}})}$ (101)
the ratio of the two solutions. Explicit computations show that, as in the confluent case, the ratio of the two components can be written as
$\gamma(x)={g_{+}(x)\over g_{-}(x)}=x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}(a,x)}=x^{2a}\left(1+x\left(a-\frac{16am_{1}m_{2}m_{3}m_{4}}{\left(1-4a^{2}\right)^{2}}\right)+\ldots\right)$ (102)
with ${\cal F}_{\rm inst}$ still is given by (22). The function ${\cal F}_{\rm inst}$ corresponds now to the prepotential of a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with four fundamental flavours.
5.2 Connection formulae
As in the confluent case, connection formulae for the Heun function can be easily derived from the standard hypergeometric transformation rules of $H^{p}_{\alpha}$. We introduce the two extra sets of hypergeometric functions
$\widetilde{H}^{0}_{\alpha}(y)$ $=$ $(e^{{-}i\pi}y)^{{-}\frac{1{+}\alpha}{2}m_{34}}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}{-}a{-}m_{3}{+}{\textstyle\frac{1{+}\alpha}{2}}m_{34},{\textstyle\frac{1}{2}}{+}a{-}m_{3}{+}{\textstyle\frac{1{+}\alpha}{2}}m_{34};1{+}\alpha m_{34};{\textstyle\frac{1}{y}}\right)$ (103)
$\widetilde{H}^{1}_{\alpha}(z)$ $=$ $z^{a{+}m_{3}{-}\frac{1}{2}}(1{-}z)^{\frac{1+\alpha}{2}\left(m_{1}{+}m_{2}\right)}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}{+}a{+}\alpha m_{1},{\textstyle\frac{1}{2}}{+}a{+}\alpha m_{2};1{+}\alpha\left(m_{1}+m_{2}\right);1{-}z\right)$
with $m_{34}=m_{3}-m_{4}$. We notice that both $\widetilde{H}^{0}_{\alpha}(y)$ and $\widetilde{H}^{1}_{\alpha}(z)$ are invariant under the sign flipping $a\to-a$, due to hypergeometric identities. The functions (103) are related to those introduced in (98) by the standard hypergeometric relations
$H^{0}_{\alpha}(y)=\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{H}^{0}_{\alpha^{\prime}}(y)\quad,\quad H^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{H}^{1}_{\alpha^{\prime}}(z)$ (104)
with
$B_{\alpha\alpha^{\prime}}$ $=$ $e^{-i\pi\left(\alpha a+m_{3}-\frac{1}{2}\right)}\frac{\Gamma(1-2\alpha a)\Gamma\left(\alpha^{\prime}\left(m_{4}-m_{3}\right)\right)}{\Gamma\left({\textstyle\frac{1}{2}}-\alpha a-\alpha^{\prime}m_{3}\right)\Gamma\left({\textstyle\frac{1}{2}}-\alpha a+\alpha^{\prime}m_{4}\right)}$
$F_{\alpha\alpha^{\prime}}$ $=$ $\frac{\Gamma(1+2\alpha a)\Gamma\left(-\alpha^{\prime}\left(m_{1}+m_{2}\right)\right)}{\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{1}\right)\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{2}\right)}$ (105)
Plugging (104) into (5.1) one finds the connection formulae
$G^{0}_{\alpha}(y)$ $=$ $\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}(y),$
$G^{1}_{\alpha}(z)$ $=$ $\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}(z)$ (106)
with
$\widetilde{G}^{0}_{\alpha}(y)$ $=$ $P_{0}(y)\widetilde{H}^{0}_{\alpha}(y){+}\widehat{P}_{0}({\textstyle\frac{x}{z}})y\widetilde{H}^{0}_{\alpha}{}^{\prime}(y)$
$\widetilde{G}^{1}_{\alpha}(z)$ $=$ $P_{1}(z)\widetilde{H}^{1}_{\alpha}(z){+}\widehat{P}_{1}(z)z\widetilde{H}^{1}_{\alpha}{}^{\prime}(z)$ (107)
describing the solution in the far region and near horizon zone. The pairs $G^{0}_{\alpha}(z)$, $G^{1}_{\alpha}(z)$, $\widetilde{G}^{0}_{\alpha}(z)$, $\widetilde{G}^{0}_{\alpha}(z)$ provide four different bases of solutions of the Heun equation. Each basis covers a different patch of the interval $z\in[0,1]$. They are related to each other by the connection formulae (106). Remarkably the Heun connection formulae (106) take the familiar hypergeometric form, with the important difference that the argument of the gamma functions depends on the non-trivial function $a(u)$. Finally, we notice that the Heun function corresponds to the minus component of the far region basis
${\rm Heun}_{-}(z)=\widetilde{G}^{0}_{-}(z)\underset{z\to 0}{\approx}1+\ldots$ (108)
Indeed in this limit $P_{0}\to 1$, $\widetilde{H}^{0}_{-}\to 1$ and (108) follows from (107).
6 Summary and conclusions
In this paper, we derive a hypergeometric representation of the confluent Heun functions $G_{0}(y)$ or $G_{1}(z)$ that resums the entire $z$ or $y$ dependence of the functions respectively, at any given order in the expansion for small $x$. In the context of gravity, these representations provide exact formulae for the wave form at given order in the PM or PN approximations respectively. The results apply to a general scattering process (along an open or closed trajectory) in the Schwarzschild geometry, where the expansion parameters $z=2M/r$ and $y=2{\rm i}\omega r$ are independent from each other. In a future publication we are going to compare this result with those coming from amplitude computations which, at a given loop order are exact in the velocities. We build four equivalent basis of solutions, $\widetilde{G}^{1}_{\alpha}$, $G^{1}_{\alpha}(z)$, $G^{0}_{\alpha}(z)$ and $\widetilde{G}^{0}_{\alpha}$ that describe with good accuracy different regions of the space time, and are connected to each other by Heun connection matrices. The final picture is summarized in the following scheme
$\begin{array}[]{lll}{\rm Near~{}horizon}&{\rm Near~{}zone}&{\rm Wave~{}zone}\\ z\approx 1&x\ll z\ll 1&z\ll x\ll 1\\ F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}&G^{1}_{\alpha}=g_{\alpha}G^{0}_{\alpha}&g_{\alpha}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}\end{array}$ (112)
where $B_{\alpha\alpha^{\prime}}$ and $F_{\alpha\alpha^{\prime}}$ are the standard hypergeometric connection matrices relating $z$ to $1/z$ or $1-z$ respectively. The whole non-triviality of the Heun connection formula is codified in the dependence of these matrices on the characteristic function $a(\ell)$. This simple result is a direct consequence of our ansatz for the solutio of the differential equation in terms of a single hypergeometric function and its derivative.
Specifying to circular orbits, we derive the wave form and energy flux at order 30PN. Given that this result is too large to appear in print we also publish ancillary mathematica files with the data needed to perform the computation. When compared against the PN expansions that one can find in literature Blanchet:2013haa ; Fujita:2012cm , these formulae resum the infinite tower of logarithmically divergent terms (tail and tail of tails), responsible for the slow convergence of the PN series, into exponentials. They also summarise in a compact form the dependence on $\ell$ of the waveforms, all the so called tail and tail of tail terms of the multipole expansion and have no logarithmic terms (responsible for the slow convergence of the PN series). The results for the waveform have been compared against those obtained using the numerical package BHToolkit, finding an agreement from 9 to 12 digits in the whole range of parameters allowed for bounded orbits.
Similar results can be also obtained for solutions of the non-confluent Heun equation and can therefore be applied to other physical problem governed by a Heun differential equation or a confluence of it. In particular, they can be easily adapted to the case of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes, cosmological perturbations Bianchi:2021xpr ; Bianchi:2021mft ; DiRusso:2024hmd ; Bianchi:2024vmi ; Bianchi:2024mlq ; Bianchi:2024rod . It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^8] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to also come back to this issue in a future publication.
Acknowledgements
F. F. and J. F. M. were partially supported by the MIUR PRIN contract 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity” and the INFN project ST&FI “String Theory and Fundamental Interactions”. Moreover they thank M. Bianchi, M. Billò, D. Bini, G. Bonelli, M.L. Frau, A.Lerda, A.Nagar, M.Panzeri, A. Tanzini for fruitful scientific exchanges. A. C. would like to thank the Department of Theoretical Physics at CERN for kind hospitality during the final stages of the present work. The research of R. P. and supported by the Armenian SCS grants 21AG-1C060 and 24WS-1C031. H. P. received support from the Armenian SCS grants 21AG-1C062 and 24WS-1C031.